OpenStudy (anonymous):

What is the percentage yield of dry polymer you obtained? Assume the polymer consists of 72% acrylate and 28% acryclic acid, with an average monomer molecular mass of 87.9g/mol. This question is from making a superabsorbent polymer through free-radical polymerization using acryclic acid, MBA, and VA-044. Would you please tell me how to calculate it asap?

OpenStudy (anonymous):

My instructor said to ignore the mass of MBA and VA-044. And the mass of the product I obtained was 0.112g.

OpenStudy (abb0t):

Assume that you have 100%, since 0.72 + 0.28 = 1 and 1 x 100 = 100%. So, that means you have 0.72 grams of acrylate, and 0.28 grams of acrylic acid. Now, to find percent yield, you have to use: \(\sf \color{red}{\frac{actual ~yield}{theoretical~ yield}}\times 100\) this gives you percent yield. What this tells you is how how efficient the reaction is under certain conditions.